8.
Use the following equation to answer the following question.
y = x^2 - 2x - 15
What is the equation of the axis of symmetry?
A. x = -1
B. x = 1
C. y = -1
D. y = 1

Answer by waynest(281)

y = x^2 - 2x - 15
y = x^2 - 2x + (2/2)^2 - (2/2)^2 - 15
y = x^2 - 2x + 1 - 1 - 15
y = (x -1)^2 - 16
axis of symmetry: x = 1